12 ms·
Because check mate is a very small subset of possible moves at the end game, I'm guessing the vast majority of games will end(?) with 2 kings moving around rand
by acadien 12y ago
Because check mate is a very small subset of possible moves at the end game, I'm guessing the vast majority of games will end(?) with 2 kings moving around randomly for all of time.
This assumes most games will make it past the hump of mid game where its possible the king's motion will be limited and a checkmate can erroneously happen, I suspect this is a rare case as well.
On a side note I wonder what kind of useful information could be mined from a huge set of random-random games. Maybe the relative power of each piece in terms of the number of pieces it captures on average? Also with some heuristic tweeks you could probably start exploring the power of various openings. Of course then its no longer random-random.
- afro88 12y agoIf it's impossible to checkmate then it's a stalemate. So if one side has one king and the other side one king (or even one king and a pawn, bishop, knight or rook) then it's a stalemate and the game ends. edit: forgot you can checkmate with just a king and a queen
- cozzyd 12y agoPawn wouldn't be stalemate because it can be promoted to a queen. I think it's also possible to checkmate with a rook + king (if you corner the king and have your king cover the escape route from check).
- umanwizard 12y agoYup, you're right! The algorithm works like this: (each "move" is an implicit "continue" in the pseudocode) while no checkmate: If your rook is about to be taken, move it to the other end of the current rank. If the enemy king is on rank n and your rook is anywhere but rank n-1, move your rook to rank n-1 (unless it can then be taken, in which case move it to the end of its current rank.) If your king is anywhere but rank n-2, move towards there. If your kings are on the same file, advance the rook to rank n (unless it can be taken...) If |file of your king - file of opponents king| is odd, make a wasted move with the rook (keeping it on the same rank). If you've gotten this far, move your king towards the opponent's king (staying on the same rank n-2).
- deleted 12y ago[deleted]
- Someone 12y agoRook + king is a simple checkmate (http://www.chesscorner.com/tutorial/basic/r_k_mate/r_k_mate.htm http://www.chesscorner.com/tutorial/basic/r_k_mate/r_k_mate....)
- stolio 12y agoYou can easily checkmate with a king and a rook, it just has to be along an edge of the board. In rare circumstances you can checkmate with a king and a single bishop: http://www.chess.com/forum/view/more-puzzles/mate-with-one-bishop http://www.chess.com/forum/view/more-puzzles/mate-with-one-b...
- lmm 12y agoAny piece (well, except the king) can checkmate with cooperation from the opponent's pieces. The standard sets (queen, rook, two bishops, bishop and knight) are those with which you can force a checkmate against an opponent who just has a king.
- dfan 12y agoStalemate is a very specific term that means that the side to move has no legal moves. If checkmate is impossible this is generally called a "draw by insufficient material".
- TillE 12y agoI saw one game quickly become a lone white king against six black pieces. The king eventually captured three of them, and then the game gave up after some number of fruitless moves. Even with that advantage, black couldn't randomly stumble across checkmate.
- joe_the_user 12y agoAre you sure you didn't end with a stalemate? That's what I got and I think once you get to the end, a stalemate with an imbalance of pieces on one side may be the most likely.
- Carrok 12y agoI just got a checkmate first go, rather quickly actually.
- mhomde 12y agoif anyone doesn't capture anything or move a pawn in 50 moves its an draw (better train on that bishop & knight endgame :)... if anyone claims it, so uh, not sure how that counts :) However if an checkmate is impossible with the remaining pieces it's automatic edit: missing pawn from sentence
- tzs 12y ago> if anyone doesn't capture or move anything in 50 moves its an draw You didn't quite state that right--every move something moves, so it is not possible to go 50 moves without moving anything. :-) The rule is 50 moves without a pawn moving or anything being captured. Also, it can be a draw, but it does not have to be. When the last 50 moves have been without a pawn move or a capture, the player on the move can claim a draw if he wishes, but does not have to do so. Recently, FIDE added a 75 move rule. It is similar to the 50 move rule, except that it is a draw even if both players wish to continue the game. (They also did a similar thing with the 3-fold repetition rule, which like the 50 move rule is only a draw if a player claims it. There is now a 5-fold repetition rule that draws even if neither player wants to claim a draw).
- randomnumber53 12y agoAlso, the reason it is 50 moves <em>without a pawn move or capture</em> is because both pawn moves and captures are irreversible. In fact, the only irreversible move that doesn't reset the 50 move rule is castling (or losing the right to castle), as having the ability to castle technically changes the position.
- breuleux 12y ago> I'm guessing the vast majority of games will end(?) with 2 kings moving around randomly for all of time. Hah, that's what I got on my second run: http://i.imgur.com/joLf3BX.png http://i.imgur.com/joLf3BX.png The game stopped right there, though. Edit: Hah! Again! http://i.imgur.com/2ItU89S.png http://i.imgur.com/2ItU89S.png
- acadien 12y agoYes, its because no other piece can take a king but a king can take all other pieces. The case for taking a king is actually quite rare!
- umanwizard 12y agoThis is a drawn game according to the rules of chess, since checkmate is impossible. Other draw conditions include: * Stalemate (=The side to move has no legal moves) * 50 moves in a row without any permanent changes happening ("permanent changes" means a pawn moving or a piece being captured) * Exactly the same game state occurs for the third time ("game state" is whose turn it is + status of castling ability + status of en passant possiblity + which pieces are where) * Agreed draw Fun fact: In normal time controls (not blitz) at competitive levels, chess games virtually never end in one of the termination conditions. I don't have statistics on hand but I would expect resignations and agreed draws account for 98%+ of grandmaster-level games.
- bo1024 12y agoAs I understand it, the threefold repetition rule does not mandate a draw, it just allows either player to claim a draw upon occurrence.
- narag 12y agoI believe 50 moves rules is the same. But in that case, one player is more interested in the draw than the other. In the case of repetition, both players are avoiding a move that gives the other the advantage. It's not so much a mandatory or not question, but if one of them is more willing to try a different move than to draw.
- hibikir 12y agoIt's not just random vs random. Two extremely weak players will take hours to finish a won game, because the player with the major material advantage does not know how to win! When teaching kids, it's not uncommon to, after teaching them how to move all the pieces, spend time teaching them how to perform the most basic of mates: KQ vs K, KRR vs K, KR vs K. Only after those three are mastered we have a good opportunity of having beginner games ending in anything other than a stalemate.
- omegaham 12y agoI still don't know how to checkmate with two bishops. I mean, I've read how to do it, but if you stuck me in an endgame situation, I'd probably draw. Luckily, I have never actually had an endgame with a king and two bishops versus a king. Same thing with king, knight, and bishop.
- UXDork 12y agoYou have to move the king into a corner and block his way out with your king.
- jobigoud 12y agoWhich is also a good recipe for a stalemate.
- mathrawka 12y agoIt's always about creating a "box" where the king cannot get out of, and make the box smaller as you force the king onto the edge or corner, where you can finally get the mate move.
- donmb 12y agoMore interesting is checkmate with two knights which is only possible if the other king makes a mistake. You cannot force it so if two good players end up in this situation, they mostly agree on a draw.
- randomnumber53 12y agoI only watched one game, but it seems that: 1) Yes, most games will end in stalemate. 2) Most wins would occur when one side has enough passed paws such that they chance upon promoting to a few queens. 3) The random-promotion significantly lowers the chances for checkmate. Promotion to a queen is so common that many players who play online chess set paws to automatically promote to queens, so they don't waste the 2 or so seconds it takes to click the "Queen" button.
- logfromblammo 12y agoThere are only a few circumstances where you would not promote to queen. The first is if promoting to queen would cause stalemate. In that case, you would probably promote to bishop instead, because rook would cause the same problem as queen, and knight is harder to force mate with. In the rare case where a bishop would cause stalemate, you promote to rook instead. The second is if promoting to knight would force checkmate faster or more surely than queen, which is so rare you practically need to play with the intent of making it happen. Either one is trivial to determine in a chess-playing program. Check 1: would a knight promotion cause immediate checkmate? If yes, then promote to knight. Check 2: would a queen promotion cause immediate stalemate? If no, then promote to queen. Check 3: would a rook promotion cause immediate stalemate? If no, then promote to rook. Check 4: would a bishop promotion cause immediate stalemate? If no, then promote to bishop. Check 5: would a knight promotion cause immediate stalemate? If no, then promote to knight. Otherwise, promote to queen and take the stalemate. Or you could simply make queen promotion the automatic default unless you advanced the pawn using a context menu to promote to a different piece.
- randomnumber53 12y agoThis comment is as far from true as the decision is far from trivial. That is to say, very far. By far the second most common piece to promote to is a knight, simply because a knight moves in a way a queen cannot. In my 10 years of playing chess, I have probably seen about 5-10 tournament games where promoting to a knight is the correct choice. Here is a FEN of an example where the only saving move for Black is to play 1...e1=N+: "8/8/8/8/8/3K4/R3p3/3k4 b - - 0 1" You are right in that the only reason to ever promote to a bishop or rook is to avoid stalemate, yet there is only one practical example of such, here: "8/2P5/8/8/3r4/8/2K5/k7 w - - 0 1". White must promote to a rook, since if they promote to a queen Black will play 1...Rc4+, forcing 2...Qxc4, which is stalemate. I cannot even find a practical example of when promoting to a bishop would the best choice.
- ekianjo 12y ago> what kind of useful information could be mined from a huge set of random-random games Whether there is a real advantage to start with white instead of black ?
- npizzolato 12y agoDoes it matter if there is an advantage to playing white vs black if both players are playing completely randomly? It should only really matter if there is an advantage when both players are playing optimally.
- ekianjo 12y agoif completely random, you may prove that there is an empirical advantage to starting the game without taking in account the level of any player, because two players playing randomly should be equivalent. I'm not entirely sure, though.
- mehrdada 12y agoThis is technically correct, but does not result in a very useful information, as P(A|B) can be a very different beast than P(A).
- ema 12y agoImagine a really simple game where there are only 2 moves: saying 1 and saying 0. The player who first says 1 wins. So of course the player who has the first turn will in practice always win, but when both players move randomly it's 50/50.
- jdoerrie 12y agoNot true. Assuming both players move completely random (both choose either 1 or 0 with 50% probability each) and independently (past actions don't influence the current event) player 1 has a 2:1 advantage, i.e. he will win with probability 2/3. This is because player 1 can win after an even number of rounds (including 0) with 50% probability. So his total chance of winning is 1/2 * (1 + 1/4 + 1/16 + 1/64 + ...) = 1/2 * 4/3 = 2/3.
- Animats 12y agoThe program stops when nobody has enough pieces to force a checkmate. In three tries, I've seen that twice, and one rook/pawn checkmate. It's a nice Javascript demo.